OpenStudy (lexi724):

MEDAL FOR CHECKING my choice is D which is 10

OpenStudy (lexi724):

and i believe im 98%right

OpenStudy (anonymous):

no

OpenStudy (anonymous):

3x = 27

OpenStudy (tkhunny):

Perhaps you could demonstrate how you managed the value for x that would make ensure the two congruent pieces are congruent?

OpenStudy (lexi724):

well A is 5 5*5-4=21 2*5+23=33 B is 6 5*6-4=26 2*6+23=35 C is 9 5*9-4=41 2*9+23=41.......................... OH

OpenStudy (lexi724):

its C

OpenStudy (anonymous):

yes

OpenStudy (lexi724):

my bad